# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v2

@@ 2.1 Review of GMLS
[…]
@@ 2.2 Approximation of tangent vectors for unknown manifolds
Once the coarse approximation to the tangent vectors is obtained, one can apply the GMLS approach to get a more accurate approximation of the manifold as well as the tangent space [37, 22, 29]. In particular, one can construct a local coordinate system with the origin $\mathbf{x}_{0}$ and the orthonormal basis $\{\bm{\tilde{t}}_{1},\bm{\tilde{t}}_{2},\bm{\tilde{n}}\}$. Then, a neighboring point $\mathbf{x}$ around $\mathbf{x}_{0}$ has the local coordinate $\left(\tilde{\theta}_{1},\tilde{\theta}_{2},\tilde{q}\right)$, where $\tilde{\theta}_{i}({\mathbf{x}})=\bm{\tilde{t}}_{i}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}})$ $(i=1,2)$ and $\tilde{q}({\mathbf{x}})=\bm{\tilde{n}}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}})$. An approximation of the manifold with higher accuracy could be calculated as follows:
@@
For a $d$ dimensional manifold $M$ embedded in $n$ dimensional ambient spaces $\mathbb{R}^{n}$ ($d\ll n$), a similar strategy can be used to approximate the $d$ tangent vectors $\{\bm{\hat{t}}_{i}\}_{i=1}^{d}$ at a given point. Specifically, we first apply the local SVD as described in above Step 1 and Step 2 to obtain a coarse approximation of the tangent vectors $\{\bm{\tilde{t}}_{i}\}_{i=1}^{d}$. Next, we define the local coordinates $\tilde{\theta}_{i}({\mathbf{x}})=\bm{\tilde{t}}_{i}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}}),1\leq i\leq d$. Rather than applying the GMLS approximation to the local normal coordinate $\tilde{q}_{i}(\mathbf{x})=\tilde{\bm{n}}_{i}(\mathbf{x}_{0})\cdot(\mathbf{x}-\mathbf{x}_{0}),1\leq i\leq n-d,$ as done in above 2D case, we instead apply the GMLS to each ambient coordinate of $\mathbf{x}-\mathbf{x}_{0}$. This approach avoided the need for generating $n-d$ normal directions $\tilde{\bm{n}}_{i}(\mathbf{x}_{0})$ and computing $n-d$ inner products in each $\tilde{q}_{i}(\mathbf{x})$ at each point, significantly reducing computational cost when $d\ll n$. A GMLS approximation of $\{\bm{\hat{t}}_{i}\}_{i=1}^{d}$ with higher accuracy can then be calculated as follows:
@@
    Solve the normal equation to obtain the regression coefficients $\{\tilde{b}_{\bm{\alpha}}^{s}\}_{1\leq\left|\bm{\alpha}\right|\leq l}$for each $s=1,\ldots,n$. The complexity is $O(nKm^{2}+nm^{3})$ at each ${\mathbf{x}}_{0}$, where $m=\left(\begin{array}[]{c}l+d\\ d\end{array}\right)$ is the number of monomial basis functions $\left\{\bm{\tilde{\theta}}^{\bm{\alpha}}|0\leq\left|\bm{\alpha}\right|\leq l\right\}$.
@@
Notably, we do not need to generate the $n-d$ normal vectors $\{\hat{\bm{n}}_{i}\}_{i=1}^{n-d}$, which would require a computational complexity of about $O(n^{3})$ when $d\ll n$. Now, the total complexity for Step 3 to Step 5 in the above algorithm is $O(n)$. We also point out that the moving least-squares based approach proposed in [55] can be used to handle point cloud data in $\mathbb{R}^{n}$.
@@ 3.1 Monge parametrization
Observe that the coordinate representation for tangent vector bases in the vicinity of ${\mathbf{x}_{0}}$ are $(1,0,\partial_{\theta_{1}}\hat{q})^{\top}$ and $(0,1,\partial_{\theta_{2}}\hat{q})^{\top}$ in the local basis $\{\bm{t}_{1},\bm{t}_{2},\bm{n}\}$ (Fig. 1(d)). For $\mathbf{x}$ close to ${\mathbf{x}_{0}}$, the tangent vector basis $\{\bm{t}_{1},\bm{t}_{2}\}\in T_{\mathbf{x}}M$ in the Cartesian coordinate system of ambient space $\mathbb{R}^{3}$ are approximated as
@@ 3.2 Approximation of the Bochner Laplacian
where $\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}$ is the $K$-nearest neighbors of $\mathbf{x}_{0}$, $\bm{u}|_{{\mathbf{x}_{0,k}}}=(u^{1}({\mathbf{x}_{0,k}}),u^{2}({\mathbf{x}_{0,k}}))^{\top}$ and $\ \mathbf{u}_{{\mathbf{x}_{0}}}=(u^{1}({\mathbf{x}_{0,1}}),u^{2}({\mathbf{x}_{0,1}}),\ldots,u^{1}({\mathbf{x}_{0,K}}),u^{2}({\mathbf{x}_{0,K}}))^{\top}\in\mathbb{R}^{2K\times 1}.$
@@ 3.2.2 Approximation of function and its derivatives in the local Monge coordinate
::=:= $\left(\bm{A}_{11}\bm{\beta}+\bm{A}_{12}\bm{\gamma}\right)\bm{t}_{1}({\mathbf{x}_{0}})+\left(\bm{A}_{21}\bm{\beta}+\bm{A}_{22}\bm{\gamma}\right)\bm{t}_{2}({\mathbf{x}_{0}}),$
where the first component of $\Delta_{B}\bm{u}(\mathbf{x}_{0})$ is $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{1}:=\bm{A}_{11}\bm{\beta}+\bm{A}_{12}\bm{\gamma}$ and the second component is $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{2}:=\bm{A}_{21}\bm{\beta}+\bm{A}_{22}\bm{\gamma}$. Here $\bm{\beta}=\left(\beta_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ and $\bm{\gamma}=\left(\gamma_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ are both $m$ by $1$ column vectors and $\{\bm{A}_{ij}\in\mathbb{R}^{1\times m}\}_{i,j=1}^{2}$ are all $1$ by $m$ row vectors. Specifically, $\bm{A}_{11}$ has three nonzero entries $24(a_{2,0})^{2}+(a_{1,1})^{2},2,$ 2 correponding to $\beta_{0,0},\beta_{2,0},\beta_{0,2}$, respectively, and $\bm{A}_{12}$ has only one nonzero entry $2a_{2,0}a_{1,1}+2a_{0,2}a_{1,1}$ correponding to $\gamma_{0,0}$. $\bm{A}_{21}$ and $\bm{A}_{22}$ are defined similarly.
@@ 3.2.3 Matrix assembly of the Bochner Laplacian
In this section, we discuss how to calculate the weights $\{\bm{w}^{k}\}$ in (13) from the weights $\{\bm{\tilde{w}}^{k}\}$ in (17) and assemble all $\{\bm{w}^{k}\}$ into the appropriate rows and columns of a sparse $2N\times 2N$ Laplacian matrix $\mathbf{L}_{B}$. So far, we have obtained the components$\ \{\bm{\tilde{u}}|_{{\mathbf{x}_{0,k}}}\in\mathbb{R}^{2\times 1}\}_{k=1}^{K}$ and the weights $\{\bm{\tilde{w}}^{k}\}_{k=1}^{K}$ to approximate $\Delta_{B}\bm{u}({\mathbf{x}_{0}})$ (see equation (17)) in the local bases $\left\{(\partial_{1}|_{\mathbf{x}_{0,k}},\partial_{2}|_{\mathbf{x}_{0,k}})\right\}_{k=1}^{K}$ of each stencil $S_{{\mathbf{x}_{0}}}$. In order to construct a sparse $2N\times 2N$ Bochner Laplacian matrix, we need to transform the weights from the local bases $\left\{\partial_{1}|_{\mathbf{x}_{0,k}},\partial_{2}|_{\mathbf{x}_{0,k}}\right\}_{k=1}^{K}$ in each $S_{{\mathbf{x}_{0}}}$ to the unified global basis $\{\bm{t}_{1}(\mathbf{x}_{0,k}),\bm{t}_{2}(\mathbf{x}_{0,k})\}_{k=1}^{K}$.
@@ 3.3.2 Numerical implementation
Note that the formulation (25) requires computing the $n-d$ normal vectors $\left\{\bm{n}_{i}({\mathbf{x}_{0}})\right\}_{i=1}^{n-d}$ at the base point which are obtained using the QR decomposition algorithm with a computational cost of $O(n^{3})$. To avoid explicitly computing these normal vectors, we instead apply the GMLS to each component of ${\mathbf{x}}=(x^{1},\ldots,x^{n})$ in the stencil $S_{{\mathbf{x}_{0}}}=\left\{{\mathbf{x}}_{0,k}\right\}_{k=1}^{K}=\{x_{0,k}^{s}\}_{k=1,\ldots,K}^{s=1,\ldots,n}$ rather than using $q^{s}$ as defined in (22). For $s=1,\ldots,n$, the $s$th component of the Taylor expansion of $\mathbf{x}(\bm{\theta})$ with respect to $\bm{\theta}=(\theta_{1},...,\theta_{d})$ at $\bm{\theta}=\bm{0}$ is given by
@@ 4 GMLS using extrinsic formulation
In this section, we approximate operators on a point cloud using the extrinsic formulation. Section 4.1 reviews the representations of the differential operators on 2D surfaces as tangential derivatives [30, 18, 14, 52, 34, 24], which could be formulated as the projection of the appropriate derivatives in the ambient space. In Section 4.2, we approximate the extrinsic formulation of the gradient and the Bochner Laplacian of vector fields using the GMLS approach [37, 15, 22, 29, 28]. The vector fields here are represented using the bases from ambient space so that the resulting discrete Bochner Laplacian is of size $3N\times 3N$ where $N$ is the number of data points and $3$ corresponds to the ambient dimension. In Section 4.3, we introduce a coordinate transformation technique to convert the $3N\times 3N$ matrix to a $2N\times 2N$ matrix where $2$ corresponds to the intrinsic dimension of the surface. In Section 4.4, we directly present the generalized results of the Bochner Laplacian for $d$-dimensional manifolds in $\mathbb{R}^{n}$. In Section 4.5, we derive the approximation formula for covariant derivatives, specifically for 2D manifolds in $\mathbb{R}^{3}$.
@@ 4.1 Review of extrinsic formulation of differential operators on manifolds
Let $T_{\mathbf{x}}M$ be the tangent space at point $\mathbf{x}\in M$ and denote a set of orthonormal tangent-vector basis by $\left\{\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})\right\}$. Then the projection matrix $\mathbf{P}$ which projects vectors in $\mathbb{R}^{3}$ to $T_{\mathbf{x}}M$ could be written as $\mathbf{P}=\sum_{i=1}^{2}\bm{t}_{i}\bm{t}_{i}^{\top}$ at any $\mathbf{x}\in M$ (see e.g. [24] for more detailed properties about $\mathbf{P}$). Subsequently, the surface gradient acting on a smooth function $f:M\rightarrow\mathbb{R}$ evaluated at $\mathbf{x}\in M$ in the Cartesian coordinates can be given by [30, 18, 14, 52, 34],
@@
where $\nabla$ is the Levi-Civita connection on $M$, ¯$\bar{\nabla}:\mathfrak{X}\left(\mathbb{R}^{n}\right)\times\mathfrak{X}\left(\mathbb{R}^{n}\right)\rightarrow\mathfrak{X}\left(\mathbb{R}^{n}\right)$ is the Euclidean connection (directional derivative) on $\mathbb{R}^{n}$ mapping $(\bm{U},\bm{Y})$ to ¯$\bar{\nabla}_{\bm{U}}\bm{Y}$, $\mathbf{P}:T\mathbb{R}^{n}\rightarrow TM$ is the orthogonal projection onto $TM$, and $\bm{U}$ and $\bm{Y}$ are extensions of $u$ and $y$ onto an open subset in $\mathbb{R}^{n}$ such that $\bm{U}|_{M}=\bm{u}$ and $\bm{Y}|_{M}=\bm{y}$, respectively. One can show that such extensions exist and the value is independent of the chosen extension and thus $\nabla_{\bm{u}}\bm{y}$ is well defined. More details about geometry can also be found in [11, 41, 8, 3] and references therein.
@@
::=:= $\left(\mathcal{H}_{1}\bm{U},\mathcal{H}_{2}\bm{U},\mathcal{H}_{3}\bm{U}\right),$
@@ 4.2 Review of local approximation of differential operators on manifolds
for any $k=1,...,K$and $s=1,2,3,$ where $\theta_{i}(\mathbf{x})$ $(i=1,2)$ are the local coordinates in (1)  and $b_{\alpha_{1},\alpha_{2}}$ are the coefficients for $(\mathcal{I}_{\mathbb{P}}\mathbf{f}_{{\mathbf{x}_{0}}})(\mathbf{x})=\sum b_{\alpha_{1},\alpha_{2}}\theta_{1}^{\alpha_{1}}(\mathbf{x})\theta_{2}^{\alpha_{2}}(\mathbf{x})$ as in (4). For each $s=1,2,3$, the above relation can be written in matrix form as,
@@
where $\mathbf{B}_{s}$ is a $K$ by $m$ matrix with $[\mathbf{B}_{s}]_{ij}=(\mathcal{G}_{s}\theta_{1}^{\alpha_{1}(j)}\theta_{2}^{\alpha_{2}(j)})(\mathbf{x}_{0,i})$, $\bm{\Phi}^{{\dagger}}$ is a $m$ by $K$ matrix defined in (4) and $\mathbf{G}_{s}=\mathbf{B}_{s}\bm{\Phi}^{{\dagger}}$ is a $K$ by $K$ matrix for approximating $\mathcal{G}_{s}$ locally in the stencil $S_{{\mathbf{x}_{0}}}$.
@@
where $\otimes$ is the Kronecker product, $\mathbf{p}_{ij}=\left(P_{ij}(\mathbf{x}_{0,1}),\ldots,P_{ij}(\mathbf{x}_{0,K})\right)^{\top}\in\mathbb{R}^{K\times 1}$, and $\delta_{kk}\in\mathbb{R}^{K\times K}$ has only one nonzero value $1$ in the $k$th row and $k$th column and has values 00 elsewhere. Note that $\mathbf{P}^{\otimes}$ is a symmetric positive definite projection matrix that maps a vector field in $\mathbb{R}^{3}$ onto the tangent space of $M$ for the $K$ neighboring points in $S_{{\mathbf{x}_{0}}}$.
@@
where $\mathbf{U}_{{\mathbf{x}_{0}}}=((\mathbf{U}_{{\mathbf{x}_{0}}}^{1})^{\top},(\mathbf{U}_{{\mathbf{x}_{0}}}^{2})^{\top},(\mathbf{U}_{{\mathbf{x}_{0}}}^{3})^{\top})^{\top}\in\mathbb{R}^{3K\times 1}$ with each component $,\mathbf{U}_{{\mathbf{x}_{0}}}^{\ell}=(U^{\ell}({\mathbf{x}_{0,1}}),$...,… , $U^{\ell}(\mathbf{x}_{0,K}))^{\top}$ $\in\mathbb{R}^{K\times 1}$ for $\ell=1,2,3$, and $\mathbf{U}_{{\mathbf{x}_{0}}}=\mathbf{P}^{\otimes}\mathbf{U}_{{\mathbf{x}_{0}}}$ since $\mathbf{U}_{{\mathbf{x}_{0}}}$ always lives in the tangent bundle of $M$.

### v3

# Generalized Moving Least-Squares Methods for Solving Vector-valued PDEs on Unknown Manifolds

@@ 2.1 Review of GMLS
$\bm{\Phi}_{kj}=\theta_{1}^{\alpha_{1}(j)}({\mathbf{x}_{0,k}})\theta_{2}^{\alpha_{2}(j)}({\mathbf{x}_{0,k}}),\quad 1\leq k\leq K,\ 1\leq j\leq m,$ (5)
@@ 2.2 Approximation of tangent vectors for unknown manifolds
[…]
@@
For a $d$ dimensional manifold $M$ embedded in $n$ dimensional ambient spaces $\mathbb{R}^{n}$ ($d\ll n$), a similar strategy can be used to approximate the $d$ tangent vectors $\{\bm{\hat{t}}_{i}\}_{i=1}^{d}$ at a given point. Specifically, we first apply the local SVD as described in above Step 1 and Step 2 to obtain a coarse approximation of the tangent vectors $\{\bm{\tilde{t}}_{i}\}_{i=1}^{d}$. Next, we define the local coordinates $\tilde{\theta}_{i}({\mathbf{x}})=\bm{\tilde{t}}_{i}({\mathbf{x}_{0}})\cdot(\mathbf{x}-{\mathbf{x}_{0}}),1\leq i\leq d$. Rather than applying the GMLS approximation to the local normal coordinate $\tilde{q}_{i}(\mathbf{x})=\tilde{\bm{n}}_{i}(\mathbf{x}_{0})\cdot(\mathbf{x}-\mathbf{x}_{0}),1\leq i\leq n-d,$ as done in above 2D case, we instead apply the GMLS to each ambient coordinate of $\mathbf{x}-\mathbf{x}_{0}$. This approach avoided the need for generating ndn-d normal directions $\tilde{\bm{n}}_{i}(\mathbf{x}_{0})$ and computing ndn-d inner products in each $\tilde{q}_{i}(\mathbf{x})$ at each point, significantly reducing computational cost when $d\ll n$. A GMLS approximation of $\{\bm{\hat{t}}_{i}\}_{i=1}^{d}$ with higher accuracy can then be calculated as follows:
@@
    Solve the normal equation to obtain the regression coefficients $\{\tilde{b}_{\bm{\alpha}}^{s}\}_{1\leq\left|\bm{\alpha}\right|\leq l}\$for each $s=1,\ldots,n$. The complexity is $O(nKm^{2}+nm^{3})$ at each ${\mathbf{x}}_{0}$, where $m=\left(\begin{array}[]{c}l+d\\ d\end{array}\right)$ is the number of monomial basis functions $\left\{\bm{\tilde{\theta}}^{\bm{\alpha}}|0\leq\left|\bm{\alpha}\right|\leq l\right\}$.
@@
Notably, we do not need to generate the ndn-d normal vectors $\{\hat{\bm{n}}_{i}\}_{i=1}^{n-d}$, which would require a computational complexity of about $O(n^{3})$ when $d\ll n$. Now, the total complexity for Step 3 to Step 5 in the above algorithm is $O(n)$. We also point out that the moving least-squares based approach proposed in [54] can be used to handle point cloud data in $\mathbb{R}^{n}$.
@@ 3.1 Monge parametrization
[…]
@@ 3.2 Approximation of the Bochner Laplacian
[…]
@@ 3.2.2 Approximation of function and its derivatives in the local Monge coordinate
$:=$ $\left(\bm{A}_{11}\bm{\beta}+\bm{A}_{12}\bm{\gamma}\right)\bm{t}_{1}({\mathbf{x}_{0}})+\left(\bm{A}_{21}\bm{\beta}+\bm{A}_{22}\bm{\gamma}\right)\bm{t}_{2}({\mathbf{x}_{0}}),$
where the first component of $\Delta_{B}\bm{u}(\mathbf{x}_{0})$ is $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{1}:=\bm{A}_{11}\bm{\beta}+\bm{A}_{12}\bm{\gamma}$ and the second component is $(\Delta_{B}\bm{u}(\mathbf{x}_{0}))_{2}:=\bm{A}_{21}\bm{\beta}+\bm{A}_{22}\bm{\gamma}$. Here $\bm{\beta}=\left(\beta_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ and $\bm{\gamma}=\left(\gamma_{\alpha_{1}(j),\alpha_{2}(j)}\right)_{j=1}^{m}\in\mathbb{R}^{m\times 1}$ are both $m$ by $1$ column vectors and $\{\bm{A}_{ij}\in\mathbb{R}^{1\times m}\}_{i,j=1}^{2}$ are all $1$ by $m$ row vectors. Specifically, $\bm{A}_{11}$ has three nonzero entries $4(a_{2,0})^{2}+(a_{1,1})^{2},2,2$ correponding to $\beta_{0,0},\beta_{2,0},\beta_{0,2}$, respectively, and $\bm{A}_{12}$ has only one nonzero entry $2a_{2,0}a_{1,1}+2a_{0,2}a_{1,1}$ correponding to $\gamma_{0,0}$. $\bm{A}_{21}$ and $\bm{A}_{22}$ are defined similarly.
@@ 3.2.3 Matrix assembly of the Bochner Laplacian
[…]
@@ 3.3.2 Numerical implementation
Note that the formulation (25) requires computing the ndn-d normal vectors $\left\{\bm{n}_{i}({\mathbf{x}_{0}})\right\}_{i=1}^{n-d}$ at the base point which are obtained using the QR decomposition algorithm with a computational cost of $O(n^{3})$. To avoid explicitly computing these normal vectors, we instead apply the GMLS to each component of ${\mathbf{x}}=(x^{1},\ldots,x^{n})$ in the stencil $S_{{\mathbf{x}_{0}}}=\left\{{\mathbf{x}}_{0,k}\right\}_{k=1}^{K}=\{x_{0,k}^{s}\}_{k=1,\ldots,K}^{s=1,\ldots,n}$ rather than using $q^{s}$ as defined in (22). For $s=1,\ldots,n$, the $s$th component of the Taylor expansion of $\mathbf{x}(\bm{\theta})$ with respect to $\bm{\theta}=(\theta_{1},...,\theta_{d})$ at $\bm{\theta}=\bm{0}$ is given by
@@ 4 GMLS using extrinsic formulation
In this section, we approximate operators on a point cloud using the extrinsic formulation. Section 4.1 reviews the representations of the differential operators on 2D surfaces as tangential derivatives [18, 14, 51, 33, 24], which could be formulated as the projection of the appropriate derivatives in the ambient space. In Section 4.2, we approximate the extrinsic formulation of the gradient and the Bochner Laplacian of vector fields using the GMLS approach [36, 15, 22, 29, 28]. The vector fields here are represented using the bases from ambient space so that the resulting discrete Bochner Laplacian is of size $3N\times 3N$ where $N$ is the number of data points and $3$ corresponds to the ambient dimension. In Section 4.3, we introduce a coordinate transformation technique to convert the $3N\times 3N$ matrix to a $2N\times 2N$ matrix where $2$ corresponds to the intrinsic dimension of the surface. In Section 4.4, we directly present the generalized results of the Bochner Laplacian for $d$-dimensional manifolds in $\mathbb{R}^{n}$. In Section 4.5, we derive the approximation formula for covariant derivatives, specifically for 2D manifolds in $\mathbb{R}^{3}$.
@@ 4.1 Review of extrinsic formulation of differential operators on manifolds
Let $T_{\mathbf{x}}M$ be the tangent space at point $\mathbf{x}\in M$ and denote a set of orthonormal tangent-vector basis by $\left\{\bm{t}_{1}(\mathbf{x}),\bm{t}_{2}(\mathbf{x})\right\}$. Then the projection matrix $\mathbf{P}$ which projects vectors in $\mathbb{R}^{3}$ to $T_{\mathbf{x}}M$ could be written as $\mathbf{P}=\sum_{i=1}^{2}\bm{t}_{i}\bm{t}_{i}^{\top}$ at any $\mathbf{x}\in M$ (see e.g. [24] for more detailed properties about $\mathbf{P}$). Subsequently, the surface gradient acting on a smooth function $f:M\rightarrow\mathbb{R}$ evaluated at $\mathbf{x}\in M$ in the Cartesian coordinates can be given by [18, 14, 51, 33],
@@
where $\nabla$ is the Levi-Civita connection on $M$, $\bar{\nabla}:\mathfrak{X}\left(\mathbb{R}^{n}\right)\times\mathfrak{X}\left(\mathbb{R}^{n}\right)\rightarrow\mathfrak{X}\left(\mathbb{R}^{n}\right)$ is the Euclidean connection (directional derivative) on $\mathbb{R}^{n}$ mapping $(\bm{U},\bm{Y})$ to $\bar{\nabla}_{\bm{U}}\bm{Y}$, $\mathbf{P}:T\mathbb{R}^{n}\rightarrow TM$ is the orthogonal projection onto $TM$, and $\bm{U}$ and $\bm{Y}$ are extensions of $u$ and $y$ onto an open subset in $\mathbb{R}^{n}$ such that $\bm{U}|_{M}=\bm{u}$ and $\bm{Y}|_{M}=\bm{y}$, respectively. One can show that such extensions exist and the value is independent of the chosen extension and thus $\nabla_{\bm{u}}\bm{y}$ is well defined. More details about geometry can also be found in [11, 40, 8, 3] and references therein.
@@
$:=$ $\left(\mathcal{H}_{1}\bm{U},\mathcal{H}_{2}\bm{U},\mathcal{H}_{3}\bm{U}\right),$
@@ 4.2 Review of local approximation of differential operators on manifolds
for any k=1,,Kk=1,...,K\ and $s=1,2,3,$ where $\theta_{i}(\mathbf{x})$ $(i=1,2)$ are the local coordinates in (1) and $b_{\alpha_{1},\alpha_{2}}$ are the coefficients for $(\mathcal{I}_{\mathbb{P}}\mathbf{f}_{{\mathbf{x}_{0}}})(\mathbf{x})=\sum b_{\alpha_{1},\alpha_{2}}\theta_{1}^{\alpha_{1}}(\mathbf{x})\theta_{2}^{\alpha_{2}}(\mathbf{x})$ as in (4). For each $s=1,2,3$, the above relation can be written in matrix form as,
@@
where $\mathbf{B}_{s}$ is a $K$ by $m$ matrix with $\mathbf{B}_{s}]_{ij}=(\mathcal{G}_{s}\theta_{1}^{\alpha_{1}(j)}\theta_{2}^{\alpha_{2}(j)})(\mathbf{x}_{0,i})$, $\bm{\Phi}^{{\dagger}}$ is a $m$ by $K$ matrix defined in (4) and $\mathbf{G}_{s}=\mathbf{B}_{s}\bm{\Phi}^{{\dagger}}$ is a $K$ by $K$ matrix for approximating $\mathcal{G}_{s}$ locally in the stencil $S_{{\mathbf{x}_{0}}}$.
@@
where $\otimes$ is the Kronecker product, $\mathbf{p}_{ij}=\left(P_{ij}(\mathbf{x}_{0,1}),\ldots,P_{ij}(\mathbf{x}_{0,K})\right)^{\top}\in\mathbb{R}^{K\times 1}$, and $\delta_{kk}\in\mathbb{R}^{K\times K}$ has only one nonzero value $1$ in the $k$th row and $k$th column and has values 0 elsewhere. Note that $\mathbf{P}^{\otimes}$ is a symmetric positive definite projection matrix that maps a vector field in $\mathbb{R}^{3}$ onto the tangent space of $M$ for the $K$ neighboring points in $S_{{\mathbf{x}_{0}}}$.
@@
where $\mathbf{U}_{{\mathbf{x}_{0}}}=((\mathbf{U}_{{\mathbf{x}_{0}}}^{1})^{\top},(\mathbf{U}_{{\mathbf{x}_{0}}}^{2})^{\top},(\mathbf{U}_{{\mathbf{x}_{0}}}^{3})^{\top})^{\top}\in\mathbb{R}^{3K\times 1}$ with each component $\mathbf{U}_{{\mathbf{x}_{0}}}^{\ell}=(U^{\ell}({\mathbf{x}_{0,1}}),$ ,..., $U^{\ell}(\mathbf{x}_{0,K}))^{\top}$ $\in\mathbb{R}^{K\times 1}$ for $\ell=1,2,3$, and $\mathbf{U}_{{\mathbf{x}_{0}}}=\mathbf{P}^{\otimes}\mathbf{U}_{{\mathbf{x}_{0}}}$ since $\mathbf{U}_{{\mathbf{x}_{0}}}$ always lives in the tangent bundle of $M$.